OpenStudy (my4babiesandme):

solve x^2+2x-4=0 i think i have the answer but want to b certain

OpenStudy (anonymous):

\[x=-\sqrt{5}-1\] or \[x=\sqrt{5}-1\] Do you need help with how to find this?

OpenStudy (my4babiesandme):

i have it as 2 ans -2 can you show how you got your answer?

OpenStudy (anonymous):

4-4*1*(-4) = 20 ==> x=[ -2+ or (-) sqrt(20)]/2 ==> x= [-2+2sqrt(5)]/2 or x= [-2-2sqrt(5)]/2

OpenStudy (anonymous):

hence, x = -1 +sqrt(5)/2 or x =-1-sqrt(5)/2

OpenStudy (my4babiesandme):

the anwer is not a square root

myininaya (myininaya):

\[x^2+2x-4=0\] add 4 on both sides \[x^2+2x=4\] add 1 on both sides \[x^2+2x+1=4+1\] we added one so that we could complete the square on the left hand side \[(x+1)^2=5\] now take square root of both sides \[x+1=\pm \sqrt{5}\] now subtract one on both sides \[x=-1 \pm \sqrt{5}\]